# When Respondents Don’t Care Anymore: Identifying the Onset of Careless Responding

### v3

@@ Abstract
Questionnaires in the behavioral and organizational sciences tend to be lengthy. However, literature suggests that survey length is a contributing factor to careless responding, with longer questionnaires yielding higher probability that participants start responding carelessly. Consequently, in long surveys a large number of participants may engage in careless responding, posing a major threat to internal validity. We propose a novel method for identifying the onset of careless responding (or an absence thereof) that searches for a changepoint in combined measurements of multiple dimensions in which carelessness may manifest, such as inconsistency and invariability. It is highly flexible, based on machine learning, and provides statistical guarantees for controlling the false positive rate. In simulation experiments, the proposed method achieves high accuracy in identifying carelessness onset and discriminates well between attentive and various types of careless responding, even when a large number of careless respondents are present. An empirical application highlights how identifying partial carelessness uncovers novel insights on careless responding behavior. Furthermore, we provide the freely available open source software package carelessonset to facilitate adoption by empirical researchers.
@@ Introduction
Research in the behavioral and organizational sciences often involves the administration of lengthy (online) self-report questionnaires. For instance, common personality measures tend to consist of hundreds of items, such as the Revised NEO Personality Inventory (Costa & McCrae,, 1992) with 240 items or the Minnesota Multiphasic Personality Inventory-2 Restructured Form (Ben-Porath & Tellegen,, 2008) with 338 items. Even if one does not use such extensive measures, the number of items can easily reach three digits by including several shorter measures. However, it is well known that questionnaire length can have a concerning adverse effect on measurement accuracy: Respondents may experience fatigue or boredom as they progress through a lengthy questionnaire, which can provoke careless responding (e.g., Bowling et al.,, 2021; Gibson & Bowling,, 2020; Galesic & Bosnjak,, 2009), in particular in online administrations (e.g., Johnson,, 2005; Meade & Craig,, 2012). We call such participants partially careless respondents as they initially provide accurate responses, but resort to careless responding after a certain item and remain careless for the remainder of the questionnaire.
@@
According to the definition of Ward & Meade, (2023), “careless responding occurs when participants are not basing their response on the item content, and it can occur when a respondent does not read an item, does not understand an item, or is unmotivated to think about what the item is asking”.11endnote: 1Perhaps more commonly used, Huang et al., (2012) define careless responding as “a response set in which the respondent answers a survey measure with low or little motivation to comply with survey instructions, correctly interpret item content, and provide accurate responses”. We prefer the definition of Ward & Meade, (2023), as it highlights that careless responding may not be intentional (cf. Huang et al.,, 2012; Ward & Pond,, 2015; Ward et al.,, 2017). Also note that Huang et al., (2012) refer to careless responding as insufficient effort responding in their definition. Other synonyms for careless responding are participant inattention (Maniaci & Rogge,, 2014), inconsistent responding (Greene,, 1978), protocol invalidity (Johnson,, 2005), and random responding (e.g., Beach,, 1989; Berry et al.,, 1992; Emons,, 2008). Notably, Schroeders et al., (2022) point out that the latter—random responding—may be a misnomer since careless responding can also arise in non-random patterns (e.g., a recurring sequence of 1-2-3-4-5). Careless responding has been identified as a major threat to the validity of research findings through a variety of psychometric and statistical issues such as lower scale reliability, errors in statistical hypothesis testing, and deteriorated model fit (e.g., Arias et al.,, 2020; Huang et al.,, 2012, 2015; Kam & Meyer,, 2015). It is widely prevalent (e.g., Bowling et al.,, 2016; Curran,, 2016; DeSimone et al.,, 2015; Meade & Craig,, 2012), suspected to be present in all questionnaire data (Ward & Meade,, 2023), and already incidence rates as low as 5% can be problematic (Arias et al.,, 2020; Credé,, 2010; Welz et al., 2024a, ; Welz et al., 2024b, ). Often viewed as merely an insidious source of measurement error, careless responding can be a valuable source of information about study participants and the study itself. For instance, carelessness has been linked to participant personality (e.g., Bowling et al.,, 2016; DeSimone et al.,, 2020) and design factors such as study length (e.g., Bowling et al.,, 2021) and item wording. In particular negatively-worded items are prone to misresponses (e.g., Swain et al.,, 2008; Weijters & Baumgartner,, 2012) triggered by carelessness. Consequently, it is widely recommended to carefully screen survey data for the presence of careless responding (e.g., DeSimone et al.,, 2015; Huang et al.,, 2015).
@@
Numerous methods for identifying careless respondents after questionnaire administration (so-called screeners) have been proposed, for instance inconsistency indicators such as psychometric synonyms or antonyms (Meade & Craig,, 2012), invariability indicators such as the longstring index (Johnson,, 2005), or multivariate outlier analyses (e.g., Curran,, 2016). We refer to Arthur et al., (2021), DeSimone et al., (2015), and Ward & Meade, (2023) for detailed reviews.22endnote: 2A popular pre-administration method for the detection of careless responding is the inclusion of detection items (e.g., Table 1 in Arthur et al.,, 2021). Such detection items are based on the presumption that an attentive respondent will respond in a specific way, while careless respondent may fail to do so. For instance, it is expected that an attentive respondent would strongly disagree to so-called bogus items such as “I am paid biweekly by leprechauns” (Meade & Craig,, 2012). Alternative types of detection items are instructed items and self-report items (Arthur et al.,, 2021). More recently, Alfons & Welz, (2024), Ozaki, (2024), and Schroeders et al., (2022) have studied machine learning techniques for identifying careless respondents. Existing methods and their implementations in freely available software packages such as careless (Yentes & Wilhelm,, 2023) make screening for careless respondents a simple task for empirical researchers, thus elevating reliability and validity of research findings and playing an instrumental role in gaining insights into the phenomenon of careless responding. In addition, many popular screeners are intuitive and easy to understand. It should be noted that common indicators typically compute a single score for each participant, with respondents deemed careless and removed from subsequent analysis if their score exceeds a certain cutoff value. Although there exist many comparative studies on such screeners (e.g., Curran,, 2016; DeSimone & Harms,, 2018; Goldammer et al.,, 2020; Huang et al.,, 2012), their behavior under partially careless responding is unfortunately not well-understood. Furthermore, excluding partially careless respondents—particularly in lengthy surveys—may leave a lot of valuable information unused, namely the respective parts of the survey in which they responded attentively.
@@
We therefore aim to address a more nuanced question: Can we identify at which point in the survey participants start responding carelessly, if at all? Indeed, there is evidence that many—though not all—careless respondents tend to begin a survey as attentive and truthful respondents, but may resort to careless responding as the survey progresses, for instance due to fatigue or boredom (Bowling et al.,, 2021; Brower,, 2018; Galesic & Bosnjak,, 2009; Gibson & Bowling,, 2020; Ward et al.,, 2017); cf. also the literature on back random responding (e.g., Clark et al.,, 2003) or partial random responding (e.g. Berry et al.,, 1992; Gallen & Berry,, 1997). Successful identification of the onset of careless responding can therefore yield valuable, complementary insights to those from existing screeners. For subsequent analysis, the issue of partially careless responding may be transformed into a missing data problem. Rather than excluding partially careless respondents altogether, researchers can restrict their primary analyses to the parts of attentive responses, treating the careless responses as missing values. Common guidelines for handling missing data (e.g., Newman,, 2014) can thereby be followed. In addition, accurately identifying the onset of careless responding may uncover implications for survey design (e.g., if many participants start responding carelessly at a similar point in the survey, it may be beneficial to rephrase or replace those items, or to place an intervention) and may lead to new theory and a deeper understanding of careless responding behavior (e.g., potentially linking certain personality traits to a tendency for an earlier or later onset of careless responding; cf. Gibbert et al.,, 2021, for theory-building based on outlying data).
@@
For the purpose of identifying at which point in the survey a given participant becomes careless, if at all, we introduce a novel method called careless onset detection in extensive rating-scale surveys (CODERS). Specifically, CODERS estimates the item at which carelessness onsets in each questionnaire participant, or an absence thereof. An important issue is that careless responding may manifest in multiple distinctive ways (e.g., Curran,, 2016; DeSimone et al.,, 2015; DeSimone & Harms,, 2018; Edwards,, 2019; Ward & Meade,, 2023). To capture various types of carelessness, CODERS can combine multiple dimensions of evidence for careless responding to construct a score that, for each item, measures if a given respondent has started responding carelessly by that item. It achieves this by searching for changepoints along a multidimensional series of indices that are potentially indicative of carelessness, such as measures of response inconsistency or invariability.
@@
We argue that the notion of a changepoint is intuitive when studying partial carelessness: if a participant starts responding carelessly, we expect a structural break in their responses. In particular, such a respondent may abandon content-based responding and resort to careless response types from the changepoint onward. CODERS is highly flexible as it does not assume a statistical model, nor does it require predefining what types of careless response types may exist, and it is primarily intended for lengthy multi-scale surveys. Moreover, CODERS builds on nonparametric estimation theory for cutoff values with statistical guarantees regarding the false positive rate (i.e., the probability of attentive respondents being incorrectly flagged as partially careless).
To the best of our knowledge, CODERS is the first attempt to systematically detect the onset of careless responding. Through extensive simulation experiments, we find that CODERS accurately estimates carelessness onset (or a lack thereof) and achieves excellent sensitivity and specificity in detecting partially careless responding. Crucially, CODERS should not be viewed as a replacement of existing screeners for careless respondents, but as a complementary tool for a related but more nuanced task, as highlighted by an empirical application. Re-examining data from a a seminal study on careless responding in Big Five personality measurements (Johnson,, 2005), we demonstrate novel insights that can be obtained by estimating onset items of partially careless responding. Reflecting on those insights, we subsequently discuss limitations and potential extensions of CODERS. Our concluding discussion highlights how CODERS complements various existing approaches for the detection of careless respondents. Finally, we provide the freely available software package carelessonset (Welz & Alfons,, 2024), which implements CODERS in the statistical programming environment R (R Core Team,, 2024) and can be downloaded from https://github.com/mwelz/carelessonset. As such, it is a useful and accessible tool for any researcher who is concerned with survey fatigue or careless responding in general.
@@
The literature generally identifies three distinct ways in which careless responding may manifest: low internal consistency of the given responses (inconsistency), low variability of given responses (invariability), and impossibly fast responses (e.g., Curran,, 2016; DeSimone et al.,, 2015; DeSimone & Harms,, 2018; Edwards,, 2019; Ward & Meade,, 2023). We expect partially careless responding to result in a change in respondent behavior in at least one of these dimensions. CODERS therefore offers a general framework for searching for a joint changepoint in any combination of measurements of these dimensions of carelessness. While the majority of this paper is focused on inconsistency and invariability, we revisit response times in a later section. As used here, CODERS thus consists of three main components:
@@
    For each participant, a nonparametric changepoint detection test (Shao & Zhang,, 2010) is applied to a combined series of participant-item level measurements of careless responding.
@@
    As a measurement of inconsistency, reconstruction errors (RE) obtained via an auto-associative neural network (autoencoder) (Kramer,, 1992) are used. An autoencoder can be viewed as a nonlinear generalization of principal component analysis (PCA), and the (participant-item level) reconstruction errors can be interpreted as the difference between observed responses and what is expected under attentive responding.
@@
    As a measurement of invariability, we propose a novel algorithm called LongStringPattern (LSP). LSP can be viewed as a generalization of the famous longstring index of Johnson, (2005) to the participant-item level, also allowing for recurring patterns involving multiple answer categories.
Regarding the first dimension, inconsistency, the autoencoder learns response patterns that characterize attentive responding (cf. Alfons & Welz,, 2024). We expect that random-like, content-independent responses cannot be learned well and are thus poorly reconstructed by the autoencoder, leading to a changepoint in reconstruction performance. Regarding the second dimension, invariability, long sequences of identical responses or fixed response patterns are not expected in well-designed surveys (e.g., using positively and negatively coded items, randomized in some form). Hence, we expect a changepoint in response variability once a respondent commences to respond carelessly through straightlining or pattern responding behavior. To obtain a granular estimate of carelessness onset, both reconstruction errors and LSP sequences are measured on the participant-item level so that a changepoint can be detected on the item-level for each participant.
Overall, CODERS attempts to capture evidence for the onset of carelessness by combining multiple indicators that are potentially indicative of such an onset, where the different indicators may capture different manifestations of carelessness. Conversely, if a respondent never becomes careless, we do not expect a changepoint in any dimension. Combining multiple indicators is a generally recommended practice to capture various types of careless responding (Goldammer et al.,, 2020; Huang et al.,, 2012; Meade & Craig,, 2012; Ward & Meade,, 2023). We provide a detailed description of our assumptions in Appendix A and additional details on CODERS in Appendix B.
@@
Ward & Meade, (2023) define internal consistency of responses as patterns that are expected based on theoretical/logical grounds or trends in the data. For instance, items that are part of the same construct are expected to correlate strongly in most participants, provided that participants are attentive. In contrast, inconsistent careless responding “generate[s] responses that fail to meet an expected level of consistency” (Ward & Meade,, 2023). Respondents may engage in inconsistent careless responding if they attempt to conceal their carelessness, for instance by randomly choosing from all response options (Ward & Meade,, 2023). For our purposes, we consider the defining characteristic of inconsistent careless responding to be content-independent responses that are approximately randomly chosen from the response options, where the probability of being chosen may differ between response options, such as preferring extreme response options.
@@
In order to identify inconsistent careless responding, we propose to use the machine learning method of autoencoders (Kramer,, 1992), which were originally developed to filter random noise in signal processing applications. An autoencoder is a neural network (e.g., Bishop,, 2006; Goodfellow et al.,, 2016) that attempts to reconstruct its input. In other words, the output variables are equal to the input variables. The idea behind reconstructing the input data is to learn the internal structures of the data, thereby separating the signal from random noise that lacks structure. Since we consider the defining characteristic of inconsistent careless responding to be near-randomly chosen responses—which may be viewed as random noise in the context of questionnaire data—we expect autoencoders to perform well in distinguishing between attentive responses and inconsistent careless responses.
@@
To achieve the goal of learning the internal structures of a dataset, autoencoders first express the input data in terms of a low-dimensional representation before reconstructing the input data again from this low-dimensional representation (cf. Kramer,, 1992, and Chapter 14.6 in Goodfellow et al.,, 2016). The use of such a low-dimensional representation makes autoencoders appealing for questionnaire data, as questionnaires typically contain multiple items to measure each construct of interest, such as attitudes, sentiments, or personality traits. In this context, the autoencoder aims to learn the patterns between the latent constructs in the low-dimensional space. For attentive responses, randomness that can be attributed to taking multiple measurements of the latent constructs is thereby filtered out. Since the amount of randomness from these multiple measurements should be relatively low for a well-designed questionnaire, attentive responses are expected to be reconstructed well from the learned patterns in the low-dimensional space of the latent constructs. Inconsistent careless responses, on the other hand, should contain little to no structure also in this low-dimensional space, and are therefore expected to be reconstructed poorly.
@@
$\bm{x}_{1}$$\bm{x}_{2}$$\bm{x}_{3}$$\bm{x}_{4}$$\bm{\widehat{z}}_{1}$$\bm{\widehat{z}}_{2}$$\bm{\widehat{x}}_{1}$$\bm{\widehat{x}}_{2}$$\bm{\widehat{x}}_{3}$$\bm{\widehat{x}}_{4}$nput LayerMapping LayerBottleneck LayerDemapping LayerOutput Layer
@@
A schematic example of the network architecture of an autoencoder is given in Figure 1. Concretely, an autoencoder is a fully connected network whose nodes are organized in multiple layers. Each node takes a linear combination of the nodes from the previous layer and subsequently applies a so-called activation function in order to allow for nonlinear relationships. A crucial part of an autoencoder is the central layer, known as the bottleneck layer, which stores the low-dimensional representations of the input data. The network architecture of an autoencoder is symmetric around this bottleneck layer such that it can be viewed as having an encoder-decoder architecture: the encoder compresses the data into the low-dimensional representations, while the decoder reconstructs the input data again. It follows that an autoencoder can be seen as a dimension reduction technique, specifically as a nonlinear generalization of principal component analysis (PCA; see Kramer,, 1991).
Hence, the number of nodes in the bottleneck layer is an important design choice in the architecture of an autoencoder. In the context of questionnaire data, we recommend setting this number equal to the number of constructs measured in the questionnaire. For example, consider data from the Revised NEO Personality Inventory (NEO-PI-R; Costa & McCrae,, 1992), which contains 240 items and measures six so-called facets of each of the Big Five personality traits. Thus, this instrument measures a total of 30 facets, such as anxiety and modesty, so that we would set the number of nodes in the bottleneck layer to 30.
Further intuition and a thorough mathematical definition of autoencoders can be found in Appendix B.2. Although there are other design choices besides the number of nodes in the bottleneck layer (see Appendix B.2.3), our software implementation in package carelessonset (Welz & Alfons,, 2024) provides default values that we recommend researchers to use. The only input required by the user is the number of constructs measured in the survey (for the number of nodes in the bottleneck layer).
@@
where $L_{j}$ denotes the number of answer categories of item $j$. Recall that we expect the reconstruction errors $\textsf{RE}_{ij}$ to be low in case of attentive responding but high in the presence of inconsistent careless responding. Hence, if participant $i$ starts providing inconsistent careless responses at item $k\in\{1,\dots,p\}$, we expect a changepoint at position $k$ in the participant’s series of $p$ reconstruction errors, as reconstruction errors are expected to be higher from item $k$ onward. An example can be found in Figure 2, where we show simulated responses of a hypothetical participant in a 300-item questionnaire (5 answer categories per item) who starts choosing answer categories completely at random from the 172nd item, resulting in systemically larger autoencoder reconstruction errors from that item onward.
@@
Note. The simulated participant started choosing answer categories completely at random from the 172nd item onward.

### v4

@@ Abstract
Questionnaires in the behavioral sciences tend to be lengthy. However, literature suggests that survey length is a contributing factor to careless responding, with longer questionnaires yielding higher probability that participants start responding carelessly. Consequently, in long surveys a large number of participants may engage in careless responding, posing a major threat to internal validity. We propose a novel method for identifying the onset of careless responding (or an absence thereof) that searches for a changepoint in combined measurements of multiple dimensions in which carelessness may manifest, such as inconsistency and invariability. It is highly flexible, based on machine learning, and provides statistical guarantees for controlling the false positive rate. In simulation experiments, the proposed method achieves high accuracy in identifying carelessness onset and discriminates well between attentive and various types of careless responding, even when a large number of careless respondents are present. An empirical application highlights how identifying partial carelessness uncovers novel insights on careless responding behavior. Furthermore, we provide the freely available open source software package carelessonset to facilitate adoption by empirical researchers.
@@ Introduction
Research in the behavioral sciences often involves the administration of lengthy (online) self-report questionnaires. For instance, common personality measures tend to consist of hundreds of items, such as the Revised NEO Personality Inventory (Costa and McCrae, 1992) with 240 items or the Minnesota Multiphasic Personality Inventory-2 Restructured Form (Ben-Porath and Tellegen, 2008) with 338 items. Even if one does not use such extensive measures, the number of items can easily reach three digits by including several shorter measures. However, it is well known that questionnaire length can have a concerning adverse effect on measurement accuracy: Respondents may experience fatigue or boredom as they progress through a lengthy questionnaire, which can provoke careless responding (e.g., Bowling et al., 2021; Gibson and Bowling, 2020; Galesic and Bosnjak, 2009), in particular in online administrations (e.g., Johnson, 2005; Meade and Craig, 2012). We call such participants partially careless respondents as they initially provide accurate responses, but resort to careless responding after a certain item and remain careless for the remainder of the questionnaire.
@@
According to the definition of Ward and Meade (2023), “careless responding occurs when participants are not basing their response on the item content, and it can occur when a respondent does not read an item, does not understand an item, or is unmotivated to think about what the item is asking”. [^1] Perhaps more commonly used, Huang et al. (2012) define careless responding as “a response set in which the respondent answers a survey measure with low or little motivation to comply with survey instructions, correctly interpret item content, and provide accurate responses”. We prefer the definition of Ward and Meade (2023), as it highlights that careless responding may not be intentional (cf. Huang et al., 2012; Ward and Pond, 2015; Ward et al., 2017). Also note that Huang et al. (2012) refer to careless responding as insufficient effort responding in their definition. Other synonyms for careless responding are participant inattention (Maniaci and Rogge, 2014), inconsistent responding (Greene, 1978), protocol invalidity (Johnson, 2005), and random responding (e.g., Beach, 1989; Berry et al., 1992; Emons, 2008). Notably, Schroeders et al. (2022) point out that the latter—random responding—may be a misnomer since careless responding can also arise in non-random patterns (e.g., a recurring sequence of 1-2-3-4-5). Careless responding has been identified as a major threat to the validity of research findings through a variety of psychometric and statistical issues such as errors in statistical hypothesis testing, deteriorated model fit, and lower scale reliability (e.g., Stosic et al., 2024; Voss, 2024; Arias et al., 2020; Huang et al., 2012, 2015). It is widely prevalent (e.g., Bowling et al., 2016; Curran, 2016; DeSimone et al., 2015; Meade and Craig, 2012), suspected to be present in all questionnaire data (Ward and Meade, 2023), and already incidence rates as low as 5% can be problematic (Arias et al., 2020; Credé, 2010; Welz et al., 2024, 2026). Often viewed as merely an insidious source of measurement error, careless responding can be a valuable source of information about study participants and the study itself. For instance, carelessness has been linked to participant personality (e.g., Bowling et al., 2016; DeSimone et al., 2020) as well as design factors such as study length (e.g., Bowling et al., 2021) and item wording. In particular negatively-worded items are prone to misresponses (e.g., Swain et al., 2008; Weijters and Baumgartner, 2012) triggered by carelessness. Consequently, it is widely recommended to carefully screen survey data for the presence of careless responding (e.g., DeSimone et al., 2015; Huang et al., 2015).
@@
Numerous methods for identifying careless respondents after questionnaire administration (so-called screeners) have been proposed, for instance inconsistency indicators such as (resampled) personal reliabilty (Jackson, 1976; Curran, 2016; see also Goldammer et al., 2024) or psychometric synonyms/antonyms (Meade and Craig, 2012), invariability indicators such as the longstring index (Johnson, 2005), or multivariate outlier analyses (e.g., Curran, 2016). We refer to Arthur et al. (2021), DeSimone et al. (2015), and Ward and Meade (2023) for detailed reviews. [^2] A popular pre-administration method for the detection of careless responding is the inclusion of detection items (e.g., Table 1 in Arthur et al., 2021). Such detection items are based on the presumption that an attentive respondent will respond in a specific way, while careless respondent may fail to do so. For instance, it is expected that an attentive respondent would strongly disagree to so-called bogus items such as “I am paid biweekly by leprechauns” (Meade and Craig, 2012). Alternative types of detection items are instructed items and self-report items (Arthur et al., 2021). More recently, Bloy et al. (2025), Alfons and Welz (2024), Ozaki (2024), and Schroeders et al. (2022) have studied machine learning techniques for identifying careless respondents. Existing methods and their implementations in freely available software packages such as careless (Yentes and Wilhelm, 2023) make screening for careless respondents a simple task for empirical researchers, thus elevating reliability and validity of research findings and playing an instrumental role in gaining insights into the phenomenon of careless responding. In addition, many popular screeners are intuitive and easy to understand. It should be noted that common indicators typically compute a single score for each participant, with respondents deemed careless and removed from subsequent analysis if their score exceeds a certain cutoff value. Although there exist many comparative studies on such screeners (e.g., Curran, 2016; DeSimone and Harms, 2018; Goldammer et al., 2020; Huang et al., 2012), their behavior under partially careless responding is unfortunately not well-understood. Furthermore, excluding partially careless respondents—particularly in lengthy surveys—may leave a lot of valuable information unused, namely the respective parts of the survey in which they responded attentively.
@@
We therefore aim to address a more nuanced question: Can we identify at which point in the survey participants start responding carelessly, if at all? Indeed, there is evidence that many—though not all—careless respondents tend to begin a survey as attentive and truthful respondents, but may resort to careless responding as the survey progresses, for instance due to fatigue or boredom (Bowling et al., 2021; Brower, 2018; Galesic and Bosnjak, 2009; Gibson and Bowling, 2020; Ward et al., 2017); cf. also the literature on back random responding (e.g., Clark et al., 2003) or partial random responding (e.g. Berry et al., 1992; Gallen and Berry, 1997). Successful identification of the onset of careless responding can therefore yield valuable, complementary insights to those from existing screeners. For subsequent analysis, the issue of partially careless responding may be transformed into a missing data problem. Rather than excluding partially careless respondents altogether, researchers can restrict their primary analyses to the parts of attentive responses, treating the careless responses as missing values. Common guidelines for handling missing data (e.g., Newman, 2014) can thereby be followed. In addition, accurately identifying the onset of careless responding may uncover implications for survey design (e.g., if many participants start responding carelessly at a similar point in the survey, it may be beneficial to rephrase or replace those items, or to place an intervention) and may lead to new theory and a deeper understanding of careless responding behavior (e.g., potentially linking certain personality traits to a tendency for an earlier or later onset of careless responding; cf. Gibbert et al., 2021, for theory-building based on outlying data).
@@
For the purpose of identifying at which point in the survey a given participant becomes careless, if at all, we introduce a novel method called careless onset detection in extensive rating-scale surveys (CODERS). Specifically, CODERS estimates the item at which carelessness onsets in each questionnaire participant, or an absence thereof. An important issue is that careless responding may manifest in multiple distinctive ways (e.g., Curran, 2016; DeSimone et al., 2015; DeSimone and Harms, 2018; Edwards, 2019; Ward and Meade, 2023). To capture various types of carelessness, CODERS can combine multiple dimensions of evidence for careless responding to construct a score that, for each item, measures if a given respondent has started responding carelessly by that item. It achieves this by searching for changepoints along a multidimensional series of indices that are potentially indicative of carelessness, such as measures of response inconsistency or invariability.
@@
We argue that the notion of a changepoint is intuitive when studying partial carelessness: if a participant starts responding carelessly, we expect a structural break in their responses. In particular, such a respondent may abandon content-based responding and resort to careless response types from the changepoint onward. CODERS is highly flexible as it does not assume a statistical model, nor does it require predefining what types of careless response types may exist, and it is primarily intended for lengthy multi-scale surveys. Moreover, CODERS builds on nonparametric estimation theory for cutoff values with statistical guarantees regarding the false positive rate (i.e., the probability of attentive respondents being incorrectly flagged as partially careless). To the best of our knowledge, only one other approach for systematically detecting the onset of careless responding has been proposed to date: The procedure of Yu and Cheng (2019) is based on item-response-theory (IRT) modeling and is more narroly focused on random responding.
[…]
@@
The literature generally identifies three distinct ways in which careless responding may manifest: low internal consistency of the given responses (inconsistency), low variability of given responses (invariability), and impossibly fast responses (e.g., Curran, 2016; DeSimone et al., 2015; DeSimone and Harms, 2018; Edwards, 2019; Ward and Meade, 2023). We expect partially careless responding to result in a change in respondent behavior in at least one of these dimensions. CODERS therefore offers a general framework for searching for a joint changepoint in any combination of measurements of these dimensions of carelessness. While the majority of this paper is focused on inconsistency and invariability, we revisit response times in a later section. As used here, CODERS thus consists of three main components:
@@
    For each participant, a nonparametric changepoint detection test (Shao and Zhang, 2010) is applied to a combined series of participant-item level measurements of careless responding.
@@
    As a measurement of inconsistency, reconstruction errors (RE) obtained via an auto-associative neural network (autoencoder) (Kramer, 1992) are used. An autoencoder can be viewed as a nonlinear generalization of principal component analysis (PCA), and the (participant-item level) reconstruction errors can be interpreted as the difference between observed responses and what is expected under attentive responding.
@@
    As a measurement of invariability, we propose a novel algorithm called LongStringPattern (LSP). LSP can be viewed as a generalization of the famous longstring index of Johnson (2005) to the participant-item level, also allowing for recurring patterns involving multiple answer categories.
Regarding the first dimension, inconsistency, the autoencoder learns response patterns that characterize attentive responding (cf. Alfons and Welz, 2024). We expect that random-like, content-independent responses cannot be learned well and are thus poorly reconstructed by the autoencoder, leading to a changepoint in reconstruction performance. Regarding the second dimension, invariability, long sequences of identical responses or fixed response patterns are not expected in well-designed surveys (e.g., using positively and negatively coded items, randomized in some form). Hence, we expect a changepoint in response variability once a respondent commences to respond carelessly through straightlining or pattern responding behavior. To obtain a granular estimate of carelessness onset, both reconstruction errors and LSP sequences are measured on the participant-item level so that a changepoint can be detected on the item-level for each participant.
Overall, CODERS attempts to capture evidence for the onset of carelessness by combining multiple indicators that are potentially indicative of such an onset, where the different indicators may capture different manifestations of carelessness. Conversely, if a respondent never becomes careless, we do not expect a changepoint in any dimension. Combining multiple indicators is a generally recommended practice to capture various types of careless responding (Goldammer et al., 2020; Huang et al., 2012; Meade and Craig, 2012; Ward and Meade, 2023). We provide a detailed description of our assumptions in Appendix A and additional details on CODERS in Appendix B.
@@
Ward and Meade (2023) define internal consistency of responses as patterns that are expected based on theoretical/logical grounds or trends in the data. For instance, items that are part of the same construct are expected to correlate strongly in most participants, provided that participants are attentive. In contrast, inconsistent careless responding “generate[s] responses that fail to meet an expected level of consistency” (Ward and Meade, 2023). Respondents may engage in inconsistent careless responding if they attempt to conceal their carelessness, for instance by randomly choosing from all response options (Ward and Meade, 2023). For our purposes, we consider the defining characteristic of inconsistent careless responding to be content-independent responses that are approximately randomly chosen from the response options, where the probability of being chosen may differ between response options, such as preferring extreme response options.
@@
In order to identify inconsistent careless responding, we propose to use the machine learning method of autoencoders (Kramer, 1992), which were originally developed to filter random noise in signal processing applications. An autoencoder is a neural network (e.g., Bishop, 2006; Goodfellow et al., 2016) that attempts to reconstruct its input. In other words, the output variables are equal to the input variables. The idea behind reconstructing the input data is to learn the internal structures of the data, thereby separating the signal from random noise that lacks structure. Since we consider the defining characteristic of inconsistent careless responding to be near-randomly chosen responses—which may be viewed as random noise in the context of questionnaire data—we expect autoencoders to perform well in distinguishing between attentive responses and inconsistent careless responses.
@@
To achieve the goal of learning the internal structures of a dataset, autoencoders first express the input data in terms of a low-dimensional representation before reconstructing the input data again from this low-dimensional representation (cf. Kramer, 1992, and Chapter 14.6 in Goodfellow et al., 2016). The use of such a low-dimensional representation makes autoencoders appealing for questionnaire data, as questionnaires typically contain multiple items to measure each construct of interest, such as attitudes, sentiments, or personality traits. In this context, the autoencoder aims to learn the patterns between the latent constructs in the low-dimensional space. For attentive responses, randomness that can be attributed to taking multiple measurements of the latent constructs is thereby filtered out. Since the amount of randomness from these multiple measurements should be relatively low for a well-designed questionnaire, attentive responses are expected to be reconstructed well from the learned patterns in the low-dimensional space of the latent constructs. Inconsistent careless responses, on the other hand, should contain little to no structure also in this low-dimensional space, and are therefore expected to be reconstructed poorly.
@@
$\bm{x}_{1}$$\bm{x}_{2}$$\bm{x}_{3}$$\bm{x}_{4}$$\bm{\widehat{z}}_{1}$$\bm{\widehat{z}}_{2}$$\bm{\widehat{x}}_{1}$$\bm{\widehat{x}}_{2}$$\bm{\widehat{x}}_{3}$$\bm{\widehat{x}}_{4}$Input LayerMapping LayerBottleneck LayerDemapping LayerOutput Layer
@@
A schematic example of the network architecture of an autoencoder is given in Figure 1. Concretely, an autoencoder is a fully connected network whose nodes are organized in multiple layers. Each node takes a linear combination of the nodes from the previous layer and subsequently applies a so-called activation function in order to allow for nonlinear relationships. A crucial part of an autoencoder is the central layer, known as the bottleneck layer, which stores the low-dimensional representations of the input data. The network architecture of an autoencoder is symmetric around this bottleneck layer such that it can be viewed as having an encoder-decoder architecture: the encoder compresses the data into the low-dimensional representations, while the decoder reconstructs the input data again. It follows that an autoencoder can be seen as a dimension reduction technique, specifically as a nonlinear generalization of principal component analysis (PCA; see Kramer, 1991).
Hence, the number of nodes in the bottleneck layer is an important design choice in the architecture of an autoencoder. In the context of questionnaire data, we recommend setting this number equal to the number of constructs measured in the questionnaire. For example, consider data from the Revised NEO Personality Inventory (NEO-PI-R; Costa and McCrae, 1992), which contains 240 items and measures six so-called facets of each of the Big Five personality traits. Thus, this instrument measures a total of 30 facets, such as anxiety and modesty, so that we would set the number of nodes in the bottleneck layer to 30.
Further intuition and a thorough mathematical definition of autoencoders can be found in Appendix B. Although there are other design choices besides the number of nodes in the bottleneck layer, our software implementation in package carelessonset (Welz and Alfons, 2026) provides default values that we recommend researchers to use. The only input required by the user is the number of constructs measured in the survey (for the number of nodes in the bottleneck layer).
@@
where $L_{j}$ denotes the number of answer categories of item $j$. Recall that we expect the reconstruction errors $\textsf{RE}_{ij}$ to be low in case of attentive responding but high in the presence of inconsistent careless responding. Hence, if participant $i$ starts providing inconsistent careless responses at item $k\in\{1,\dots,p\}$, we expect a changepoint at position $k$ in the participant’s series of $p$ reconstruction errors, as reconstruction errors are expected to be higher from item $k$ onward. An example can be found in Figure 2, where we show simulated responses of a hypothetical participant in a 300-item questionnaire (5 answer categories per item) who starts choosing answer categories completely at random from the 183rd item, resulting in systemically larger autoencoder reconstruction errors from that item onward.
@@
Note. The simulated participant started choosing answer categories completely at random from the 183rd item onward.